A sum of $3500 is split between two investments, one paying 9% yearly interest and the other 11%. If the total yearly interest amounted to $354, how much was invested at each rate?
Answer by checkley75(3666)

.11*x+.09(3500-x)=354
.11x+315-.09x=354
.02x=354-312
.02x=39
x=39/.02
x=1950 amount of the money invested @11%.
3500-1950=1550 amount invested @ 9%.
proof
.11*1950+.09*1550=354
214.5+139.5=354
354=354